OpenStudy (anonymous):

Solve, using the substitution method. d + e = 2 d – e = 4

OpenStudy (anonymous):

use the 2nd eq so d=4+e now use this in the 1st eq 4+e+e=2 4+2e=2 2e=-2 e=-1 now use this value in either eq to get d=3

OpenStudy (anonymous):

These are My Anwser Choices: A. The solution is (3, –1). B. The solution is (5, 1). C. There is no solution. D. There are an infinite number of solutions.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Can You Tell M How You Get Tht Anwser plz?

OpenStudy (anonymous):

*Me

OpenStudy (anonymous):

yes. look at my calculations and you will see that I got 3 and -1

OpenStudy (anonymous):

Ok . Thank You .. I Appreciate Ur Help ^_^